# On the Convergence of HalpernSGD

### v1

@@ 1 Introduction
The second is the adaptive family, where learning rates are chosen coordinate-wise (or element-wise), including Adam [22], RMSProp [18], AdaGrad [14], and AdaDelta [37]. Adaptive schemes are commonly preferred in settings where rapid progress during training is crucial, e.g., GANs [17], transformers [33], and deep reinforcement learning [24]. At the same time, it is by now well documented that the convergence behavior of Adam-type methods can be subtle: counterexamples and non-convergence regimes are discussed in [28, 38], and oscillatory phenomena (including $2$-limit cycles) are reported in [6]. Moreover, in [16, 12], we comment on an imprecision in the original Adam convergence analysis, specifically in Lemma 10.3 of [22].
@@
Within the same fixed-point paradigm, the Halpern iteration is known to exhibit a faster asymptotic-regularity behavior under standard control conditions achieving the sharper rate $\mathcal{O}(n^{-1})$, and it has attracted notable interest in optimization; see, e.g., [13, 23, 36]. Motivated by this phenomenon, we consider an anchored scheme obtained by blending the Halpern iteration with the Gradient Descent (GD) step. In its deterministic form (HalpernGD), the update reads
@@
From a practical standpoint, however, full-gradient steps such as (1) are often prohibitive in large-scale learning. Mini-batch gradients provide nearly the same descent direction at a fraction of the cost, and small batches may even act as an implicit regularizer [21]. This leads to the stochastic Halpern-type iteration (HalpernSGD), which can be written as
@@
where $\{U_{n}\}$ is a martingale difference sequence known as noise and $\epsilon_{n}\downarrow 0$ is a diminishing step size. This optimizer, HalpernSGD, was proposed and experimentally investigated in our earlier work [16, 12], where we observed that the anchoring mechanism can substantially reduce the number of training epochs needed to reach a given accuracy. Additional qualitative experiments on standard nonconvex benchmarks (Rastrigin and Rosenbrock functions) further illustrate distinctive optimization trajectories and stability patterns when compared to SGD and Adam.
The purpose of this paper is to provide a rigorous convergence analysis of HalpernSGD (3) under a classical convex $L$-smooth setting and standard stochastic assumptions. Here we establish almost sure convergence of the iterates toward a distinguished minimizer characterized as the projection of the anchor $u$ onto the solution set. Our analysis relies on a suitable coupling between $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$ and on control conditions in the spirit of [35], which are natural in Halpern-type schemes. These results fill a gap in the current literature on HalpernSGD optimizer and provide a foundation for future developments, including the systematic study of anchored variants of adaptive optimizers.
Section 3 collects notation and assumptions, Section 4 contains the boundedness lemma and the almost sure convergence theorem and Section 5 concludes the paper and briefly outlines directions for future work.
@@ 2 Related Work and Contributions
The Halpern iteration originated as a fixed-point method for nonexpansive mappings and, under standard control conditions on the weight sequence, guarantees strong convergence toward the fixed point selected by the anchor, typically the fixed point closest to the anchor, see e.g. [19, 34, 35].
@@
A starting point is [10], which studies stochastic monotone inclusions and combines variants of Halpern with a recursive PAGE-type estimator to obtain variance that decreases as $O(\varepsilon^{2}/k)$, through data-driven choices of the batch-sizes. The work assumes:
  1. 1.
    unbiasedness,
  2. 2.
    bounded variance,
  3. 3.
    a multi-point oracle assumption that allows querying the oracle at multiple points using the same seed, and
  4. 4.
    a Lipschitz property in expectation.
This line therefore puts at its core the idea that, with constant-variance noise, additional mechanisms are needed to make the effective error decay [10].
@@
A related perspective, but more related to fixed-point theory, is developed in [8], which analyzes a stochastic Halpern iteration in normed spaces and discusses applications in reinforcement learning. Here, the stochastic oracle is modeled as a random approximation $\widetilde{T}(x,\xi)$ of a nonexpansive operator $T$ and the almost sure convergence to a fixed point under the undesiderable hypothesis that $\sum_{n}\|U_{n}\|<+\infty,$ where $\{U_{n}\}$ is the error sequence (Theorem 3.10).
@@
More recently, [26] proposes a general framework of stochastic Halpern scheme for nonexpansive operators, where stochasticity enters as additive perturbations $\xi_{n},\delta_{n}.$ Here, the primary objective is to establish the asymptotic regularity of the sequence of iterates, that is, the convergence $|x_{n+1}-x_{n}|\to 0$. This type of result is obtained under standard conditions on $\alpha_{n}$ and $\beta_{n}$, together with the convergence of $\sum_{n}\mathbb{E}|\delta_{n+1}-\delta_{n}|$ and of $\sum_{n}\mathbb{E}|\xi_{n+1}-\xi_{n}|$.
@@
Alongside the works specifically focused on Halpern-type schemes, general results on stochastic fixed-point iterations are also relevant. In particular, [8] analyzes Stochastic Krasnosel’skiĭ–Mann (SKM) iterations for nonexpansive operators, where the noise sequence is modeled as additive stochastic perturbations with zero mean and finite second moment, that is, $\mathbb{E}[U_{n}\mid\mathcal{F}_{n-1}]=0$ and $\mathbb{E}|U_{n}|^{2}<\infty$. Under standard stepsize assumptions, the authors establish almost sure convergence of the SKM iterates to a fixed point of $T$, without requiring the variance to be uniformly bounded.
@@
In our work we place ourselves in this latter perspective. Indeed, we study HalpernSGD as an “anchored” scheme for $L$-smooth convex minimization, modeling gradients stochastically perturbed by a unbiased noise. Almost sure convergence in our setting is established in a different way than in the works reviewed above. Specifically, we study almost sure convergence to a minimization point, that is, to a zero of the gradient of the objective function $f$, by introducing a second control sequence $\{\varepsilon_{n}\}$ that plays the role of a learning rate. This additional degree of control allows us to impose only a mild condition on the noise, namely the summability condition
$\sum_{n}\varepsilon_{n}^{2}\,\|U_{n}\|^{2}<+\infty,$
@@
which ensures that the cumulative effect of the stochastic perturbations remains tractable. As a consequence, we are able to handle almost sure convergence of the entire sequence $\{x_{n}\}$ even in the presence of noise with merely bounded variance. This, in turn, allows the Halpern framework to be applied in settings where variance reduction or vanishing-noise assumptions are not available.
@@
By exploiting a quadratic-mean recurrence and lemma proved by Xu in [35], we first prove convergence in $L^{2}$ and then almost sure convergence of the whole sequence toward the minimizer selected by the anchor, $x^{\star}=P_{S}(u)$.
@@
In summary, compared to works that require variance reduction to obtain convergence in the stochastic Halpern case [10, 8, 26], our contribution aims at an almost sure guarantee of convergence under bounded-variance assumptions and is based on an appropriate choice and coupling of two control sequences that make the cumulative effect of the noise tractable.
@@ 3 Preliminaries
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $\{\mathcal{F}_{n}\}_{n\geq 0}$ be a filtration such that $x_{n}$ (see Equation (3)) is $\mathcal{F}_{n}$-measurable. We model stochastic gradients through a martingale-difference noise sequence $\{U_{n}\}_{n\geq 0}$ satisfying
$\mathbb{E}[\,U_{n}\mid\mathcal{F}_{n}\,]=0\qquad\text{a.s.}\quad\forall n\geq 0,$ (6)
together with a bounded conditional second moment:
$\mathbb{E}\bigl[\|U_{n}\|^{2}\mid\mathcal{F}_{n}\bigr]\leq\sigma^{2}\qquad\text{a.s.}\quad\forall n\geq 0,$ (7)
for some $\sigma\geq 0$. For stepsizes $\{\epsilon_{n}\}_{n\geq 0}\subset(0,\infty)$ and Halpern weights $\{\alpha_{n}\}_{n\geq 0}\subset(0,1)$, the stochastic forward (gradient) operator is
@@
When $U_{n}\equiv 0$, (9) reduces to the deterministic Halpern iterative scheme.
In this context, a decreasing stepsize sequence $\{\epsilon_{n}\}$ is required to balance progress and noise reduction when moving from deterministic to stochastic gradients. Classical stochastic approximation theory [30] prescribes the following conditions:
@@
We also recall the standard gradient-norm bound (sometimes called a “Polyak–Łojasiewicz-type inequality” for smooth convex functions):
@@ 4 Convergence Analysis of HalpernSGD
Following the standard setting used in stochastic fixed-point theory (see, e.g.,[9]), the perturbations $U_{n}$ are assumed to satisfy the martingale difference conditions:
$\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0,\qquad\mathbb{E}[\|U_{n}\|^{2}\mid\mathcal{F}_{n}]\leq\sigma^{2}\quad\text{a.s. for all }n\in\mathbb{N},$
for some constant $\sigma>0$. The stepsizes $\{\epsilon_{n}\}$ and Halpern weights $\{\alpha_{n}\}$ are real sequences satisfying
@@
Let $x^{*}$ be a fixed point of $T$, and consequently $\nabla f(x^{*})=0$. By definition of the HalpernSGD iteration, we have
@@
    By the properties of conditional expectation,
    $\mathbb{E}\big[\mathbb{E}[\|x_{n+1}-x^{*}\|^{2}\mid\mathcal{F}_{n}]\big]=\mathbb{E}\|x_{n+1}-x^{*}\|^{2}.$
@@
We remark that $\gamma_{n}\in(0,1)$. Indeed, $\gamma_{n}=\alpha_{n}-\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, and by the coupling condition $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}$, we have $\gamma_{n}>0$. Then, iterating the inequality gives
@@
Hence, the sequence $\{a_{n}\}$ is bounded, that is, $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}<\infty.\qed$
@@
From a practical standpoint, in view of the experimental analysis, the condition $\alpha_{n}-\epsilon_{n}^{2}L^{2}>0$ can be simplified through a sufficient asymptotic requirement. Indeed, by assuming
@@
Theorem 4.2.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $L$-smooth, $x^{*}:=P_{S}(u)$ and $\{x_{n}\}$ be generated by (9). Assume the stochastic setting and the step-size conditions stated above, together with $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$.
@@
$x_{n}\xrightarrow{\text{\;a.s.}}x^{*},$
i.e., $x^{*}$ is the unique point in $S$ satisfying the variational inequality
@@
We show that $\mathbb{E}\|\nabla f(x_{n})\|^{2}\to 0$ under the stated assumptions. By definition of HalpernSGD iteration, we have
@@
$\|x_{n+1}-x^{*}\|^{2}=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},\,x_{n}-x^{*}-\epsilon_{n}g_{n}\rangle+(1-\alpha_{n})^{2}\|x_{n}-x^{*}-\epsilon_{n}g_{n}\|^{2}=$
@@
$(1-\alpha_{n})^{2}\bigl(\|x_{n}-x^{*}\|^{2}-2\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+\epsilon_{n}^{2}\|g_{n}\|^{2}\bigr).$
@@
$(1-\alpha_{n})^{2}\|x_{n}-x^{*}\|^{2}-2(1-\alpha_{n})^{2}\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\|g_{n}\|^{2}.$
@@
Next, from Lemma 4.1, the sequence of iterates is bounded in mean square, that is, there exists a constant $C>0$ such that $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}\leq C$. Thus, by applying the Cauchy–Schwarz inequality and the $L$-Lipschitz continuity of $\nabla f$, we have $|\langle\nabla f(x_{n}),u-x^{*}\rangle|\leq\|\nabla f(x_{n})\|\,\|u-x^{*}\|\leq L\sqrt{C}\,\|u-x^{*}\|.$ Hence,
$2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle\nabla f(x_{n}),u-x^{*}\rangle\leq 2\alpha_{n}(1-\alpha_{n})\epsilon_{n}L\sqrt{C}\,\|u-x^{*}\|.$
Finally, by the martingale-difference assumption, $\mathbb{E}\langle U_{n},\nabla f(x_{n})\rangle=\mathbb{E}\big[\langle\mathbb{E}[U_{n}\mid\mathcal{F}_{n}],\,\nabla f(x_{n})\rangle\big]=0,$ and then, by bounded conditional variance assumption, we have
@@
$\quad+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|.$ (17)
By the coupling $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$ we have $\epsilon_{n}\leq 1/L$ for all sufficiently large $n$, whence $\Bigl(\tfrac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)\ \geq\ \tfrac{\epsilon_{n}}{L}.$ Thus, setting $\beta_{n}:=\frac{(1-\alpha_{n})^{2}\epsilon_{n}}{L\alpha_{n}}>0$, (17) implies
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}-\mathbb{E}\|x_{n}-x^{*}\|^{2}$ $\leq\underbrace{\alpha_{n}^{2}\|u-x^{*}\|^{2}+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|}_{=:R_{n}}$
$\quad+\alpha_{n}\,\underbrace{\Bigl(2\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle-\beta_{n}\,\mathbb{E}\|\nabla f(x_{n})\|^{2}\Bigr)}_{=:\,\delta_{n}}.$
With the definitions of $R_{n}$ and $\delta_{n}$, the previous inequality becomes
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}-\mathbb{E}\|x_{n}-x^{*}\|^{2}\;\leq\;R_{n}+\alpha_{n}\,\delta_{n}.$ (18)
By the step-size assumptions, $\sum_{n}\alpha_{n}^{2}<\infty$, $\sum_{n}\epsilon_{n}^{2}<\infty$ and $\sum_{n}\alpha_{n}\epsilon_{n}<\infty$, hence $\sum_{n}R_{n}<\infty$. Summing (18) from $n=0$ to N1N-1 and using the telescoping of the left-hand side, we get
$\sum_{n=0}^{N-1}\alpha_{n}\delta_{n}\ \geq\ \mathbb{E}\|x_{N}-x^{*}\|^{2}-\mathbb{E}\|x_{0}-x^{*}\|^{2}-\sum_{n=0}^{N-1}R_{n}\ \geq\ -\,\Bigl(\mathbb{E}\|x_{0}-x^{*}\|^{2}+\sum_{n=0}^{\infty}R_{n}\Bigr)=:-\,\widetilde{M}$,
for all $N$, with some finite constant $\widetilde{M}$$\infty$ independent of $N$. Since $\sum_{n=0}^{N-1}\alpha_{n}\delta_{n}$ is bounded from below and $\sum_{n}\alpha_{n}=+\infty$, it follows that a null subsequence $\{\delta_{n_{k}}\}$ must exist, i.e.

### v2

@@ 1 Introduction
The second is the adaptive family, where learning rates are chosen coordinate-wise (or element-wise), including Adam [26], RMSProp [22], AdaGrad [17], and AdaDelta [46]. Adaptive schemes are commonly preferred in settings where rapid progress during training is crucial, e.g., GANs [21], transformers [41], and deep reinforcement learning [29]. At the same time, it is by now well documented that the convergence behavior of Adam-type methods can be subtle: counterexamples and non-convergence regimes are discussed in [35], and oscillatory phenomena (including $2$-limit cycles) are reported in [7]. Moreover, in [19, 14], we comment on an imprecision in the original Adam convergence analysis, specifically in Lemma 10.3 of [26].
@@
Within the same fixed-point paradigm, the Halpern iteration is known to exhibit a faster asymptotic-regularity behavior under standard control conditions achieving the sharper rate $\mathcal{O}(n^{-1})$, and it has attracted a particular interest in optimization; see, e.g., [16, 28, 45]. Motivated by this phenomenon, we consider an anchored scheme obtained by blending the Halpern iteration with the Gradient Descent (GD) step. In its deterministic form (HalpernGD), the update reads
@@
In [20], the authors introduced a variant of the Halpern iteration, where the step size $\eta$ is replaced by a vanishing sequence $\{\epsilon_{n}\}$, and they established strong convergence to a solution under suitable control conditions on $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$. In particular, their results imply the strong convergence of the sequence
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl(x_{n}-\epsilon_{n}\partial f(x_{n})\bigr),$ (3)
where $\partial f$ is the subdifferential of a proper, lower semicontinuous, convex function $f$ which can also be unbounded.
Here, the possibly unboundedness of the subdifferential is mitigated by the vanishing step size $\epsilon_{n}$, which prevents the iterates from diverging and allows to establish strong convergence to a minimizer.
From a practical standpoint, however, full-gradient steps such as (2) or (3) are often prohibitive in large-scale learning. Mini-batch gradients provide nearly the same descent direction at a fraction of the cost, and small batches may even act as an implicit regularizer [25]. This leads to the stochastic Halpern-type iteration (HalpernSGD), which can be written as
@@
where $\{X_{n}\}$ beacomes a sotchastic process over a probability space $\Omega$, $\{U_{n}\}$ is a martingale difference sequence and $\epsilon_{n}\downarrow 0$ is a vanishing step size.
The remainder of this paper is organized as follows. Section 3 introduces the notation, recalls fundamental properties of smooth convex functions, and states the assumptions on the stochastic perturbations and control sequences. Section 4 presents the main convergence results: Lemma 4.1 establishes mean-square boundedness of the iterates, Theorem 4.2 proves almost sure convergence of the entire sequence. Finally, Section 5 summarizes the contributions and discusses some avenues for future research, including adaptive variants and empirical benchmarking.
@@ 2 Related Work and Contributions
The Halpern iteration originated as a fixed-point method for nonexpansive mappings and, under standard control conditions on the weight sequence, guarantees strong convergence toward the anchor-selected fixed point, namely the metric projection of the anchor onto the fixed-point set, see e.g. [23, 42, 43].
Recent works have clarified sharp non-asymptotic residual guarantees for Halpern-type schemes. Early non-asymptotic $O(1/k)$ residual bounds for viscosity/Halpern-type iterations can be found e.g. in [39]. More recently, worst-case analyses have identified (near-)minimax optimal parameter choices and tight bounds for broad classes of Mann iterations; see [15, 9]. Related optimal-complexity and matching lower-bound results for fixed-point problems are developed in [32].
@@
A starting point is [12], which studies stochastic monotone inclusions and combines variants of Halpern with a recursive PAGE-type estimator to obtain a mean-square error (variance) bound for the estimator that decreases as $O(\varepsilon^{2}/k)$, via data-driven choices of the batch sizes. Anyway he work assumes a bounded variace and a multi-point oracle assumption, which allows to query the stochastic oracle at multiple points under common randomness (i.e., using the same sample/seed). property in expectation. This line therefore puts at its core the idea that, with constant-variance noise, additional mechanisms are needed to make the effective error decay [12].
@@
A related perspective, closer to fixed-point theory, is developed in [11], which analyzes a stochastic Halpern iteration in normed spaces and discusses applications in reinforcement learning. Here, the stochastic oracle is modeled as a random approximation $\widetilde{T}(x,\xi)$ of a nonexpansive operator $T$ and the almost sure convergence to a fixed point under the (rather strong) summability condition $\sum_{n}\mathbb{E}\|U_{n}\|<+\infty,$ where $\{U_{n}\}$ is the error sequence (Theorem 3.10).
@@
Related “anchoring” ideas are also leveraged to obtain near-optimal sample complexity in reinforcement learning: [27] proposes an anchored (Halpern-type) approach for computing $\varepsilon$-optimal policies in MDPs, with near-optimal sample and time complexity guarantees. In this direction, [33] proposes a general framework of stochastic Halpern scheme for nonexpansive operators, where stochasticity enters as additive perturbations $\xi_{n},\delta_{n}.$ Here, the primary objective is to establish the asymptotic regularity of the sequence of iterates, that is, the convergence $\|x_{n+1}-x_{n}\|\to 0$.
This type of result is obtained under standard control conditions on $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ (e.g., including bounded-variation requirements), together with suitable summability assumptions on the perturbation increments, such as $\sum_{n}\mathbb{E}\|\delta_{n+1}-\delta_{n}\|<+\infty$ and $\sum_{n}\mathbb{E}\|\xi_{n+1}-\xi_{n}\|<+\infty$, under a mild boundedness-in-expectation assumption on the iterates.
@@
Alongside the works specifically focused on Halpern-type schemes, general results on stochastic fixed-point iterations are also relevant. In particular, [10] analyzes Stochastic Krasnosel’skiĭ–Mann (SKM) iterations for nonexpansive operators, where the noise sequence is modeled as additive stochastic perturbations with zero mean and finite second moment, that is, $\mathbb{E}[U_{n}\mid\mathcal{F}_{n-1}]=0$ and $\mathbb{E}\|U_{n}\|^{2}<\infty$. Under standard stepsize assumptions, the authors establish almost sure convergence of the SKM iterates to a fixed point of $T$, without requiring the variance to be uniformly bounded.
@@
In our work we place ourselves in this latter perspective. Indeed, we study HalpernSGD as an “anchored” scheme for $L$-smooth convex minimization, modeling gradients stochastically perturbed by an unbiased noise sequence. Almost sure convergence in our setting is established in a different way than in the works reviewed above. Specifically, we study almost sure convergence to a minimizer, that is, to a zero of the gradient of the objective function $f$, by introducing a second control sequence $\{\epsilon_{n}\}$ that plays the role of a learning rate. This additional degree of control allows us to impose only a mild condition on the noise, namely the summability condition
$\sum_{n}\epsilon_{n}^{2}\,\|U_{n}\|^{2}<+\infty\qquad\text{a.s.},$
@@
which, under the bounded-variance assumption and $\sum_{n=0}^{\infty}\epsilon_{n}^{2}<\infty$, holds automatically almost surely and ensures that the cumulative effect of the stochastic perturbations remains tractable. As a consequence, we are able to handle almost sure convergence of the entire sequence $\{x_{n}\}$ even in the presence of noise with merely bounded variance. This, in turn, allows the Halpern framework to be applied in settings where variance reduction or vanishing-noise assumptions are not available.
@@
By exploiting a quadratic-mean recurrence and lemma proved by Xu in [43, 44], we first establish mean-square convergence, i.e., $\mathbb{E}\|x_{n}-x^{*}\|^{2}\to 0$ , and we then strengthen the result to almost sure convergence of the entire sequence to the anchor-selected minimizer $x^{*}=P_{S}(u)$.
Convergence-rate analyses for Halpern-type stabilized stochastic gradient schemes, and connections to variance growth conditions beyond the classical uniformly bounded-variance model, have recently been developed in [31, 1]; see in particular Section 3 of [1].
@@
In summary, existing stochastic analyses of Halpern-type schemes typically rely either on (i) variance reduction / increasing-batch mechanisms to enforce a decreasing effective noise level [12], or on (ii) summability/vanishing conditions on the perturbations to prevent error accumulation [11, 33]. General stochastic fixed-point results for nonexpansive maps (not necessarily Halpern-type) are also relevant; see [10]. Our contribution aims at an almost sure guarantee of convergence under bounded-variance assumptions and is based on an appropriate choice and coupling of two control sequences that make the cumulative effect of the noise tractable.
@@ 3 Preliminaries
We model stochastic gradients via a martingale-difference noise sequence $\{U_{n}\}_{n\geq 0}$, whose assumptions will be stated explicitly later on. Denoting the step sizes by $\{\epsilon_{n}\}_{n\geq 0}\subset(0,\infty)$ and the Halpern weights by $\{\alpha_{n}\}_{n\geq 0}\subset(0,1)$, we define the stochastic forward (gradient) operator as
@@
When $U_{n}\equiv 0$, (8) reduces to the deterministic Halpern iterative scheme. In this deterministic regime, strong convergence is established in [20] under the following coupled conditions
  • $\lim_{n\to\infty}\alpha_{n}=0,\quad\sum_{n=0}^{\infty}\alpha_{n}=+\infty,\quad\sum_{n=0}^{\infty}\epsilon_{n}^{\,2}<+\infty,$
  • $\lim_{n\to\infty}\dfrac{\epsilon_{n}}{\alpha_{n}}=0,\quad\sum_{n=0}^{\infty}\epsilon_{n}\alpha_{n}<+\infty.$
In the stochastic context, a decreasing stepsize sequence $\{\epsilon_{n}\}$ is required to balance progress and noise reduction when moving from deterministic to stochastic gradients. Classical stochastic approximation theory [37] prescribes the following conditions:
@@
We also recall the standard smooth-convex gradient lower bound:
@@ 4 Convergence Analysis of HalpernSGD
The stepsizes $\{\epsilon_{n}\}$ and Halpern weights $\{\alpha_{n}\}$ are real sequences satisfying
@@
Remark 1 (A concrete admissible choice of $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$).
We stress that the above coupled conditions on the Halpern weights and step sizes are not particularly restrictive. For instance, the one-parameter choice
$\alpha_{n}:=\frac{1}{n+1},\qquad\epsilon_{n}:=\frac{1}{(n+1)^{p}},\qquad p\in\Bigl(\tfrac{1}{2},1\Bigr),$
satisfies all the stated assumptions. In addition, this example also satisfies $\lim_{n\to\infty}\frac{\alpha_{n}}{\epsilon_{n}^{2}}=+\infty,$ which is a condition required at a later stage of our convergence analysis.
As for the stochastic perturbations $U_{n}$, we assume the standard martingale difference condition (see, e.g., [10])
$\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0\quad\text{a.s.}$
In place of the classical uniformly bounded-variance assumption, we impose the summability condition
$\sum_{n=0}^{\infty}\epsilon_{n}^{2}\,\mathbb{E}\!\left[\|U_{n}\|^{2}\mid\mathcal{F}_{n}\right]<+\infty\quad\text{a.s.}$
This requirement is strictly weaker than the standard bounded conditional variance hypothesis. For a deeper discussions, we refer to [1].
@@
Let $x^{*}\in S$, i.e. $\nabla f(x^{*})=0$. By definition of the HalpernSGD iteration, we have
@@
    Taking total expectations on both sides yields the corresponding inequality for $\mathbb{E}\|x_{n+1}-x^{*}\|^{2}$.
@@
We remark that $\gamma_{n}\in[0,1)$. Indeed, $\gamma_{n}=\alpha_{n}-\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, and by the coupling condition $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}$, we have $\gamma_{n}\geq 0$.
Then, iterating the inequality gives
@@
Hence, the sequence $\{a_{n}\}$ is bounded, that is, $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}<\infty.$
@@
From a practical standpoint, in view of the experimental analysis, the nonnegativity condition $\alpha_{n}-\epsilon_{n}^{2}L^{2}\geq 0$ (equivalently, $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$) can be ensured through a simple sufficient asymptotic requirement.
Indeed, by assuming
@@
Theorem 4.2 (Mean-square convergence).
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $L$-smooth, $S:=\arg\min f\neq\emptyset$, $x^{*}:=P_{S}(u)$, and let $\{x_{n}\}$ be generated by (8). Assume the stochastic setting and the step-size conditions stated above, together with
$(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1\qquad\forall n.$
@@
$\mathbb{E}\|x_{n}-x^{*}\|^{2}\longrightarrow 0.$
In particular, $x_{n}\to x^{*}$ in probability, and there exists a subsequence $(x_{n_{k}})$ such that $x_{n_{k}}\to x^{*}$ almost surely. Moreover, $x^{*}$ is the unique point in $S$ satisfying the variational inequality
@@
We show that $\mathbb{E}\|x_{n}-x^{*}\|^{2}\to 0$ under the stated assumptions. By definition of HalpernSGD iteration, we have
@@
$\|x_{n+1}-x^{*}\|^{2}=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},\,x_{n}-x^{*}-\epsilon_{n}g_{n}\rangle+(1-\alpha_{n})^{2}\|x_{n}-x^{*}-\epsilon_{n}g_{n}\|^{2}$
@@
$\qquad+(1-\alpha_{n})^{2}\bigl(\|x_{n}-x^{*}\|^{2}-2\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+\epsilon_{n}^{2}\|g_{n}\|^{2}\bigr).$
@@
$\qquad+(1-\alpha_{n})^{2}\|x_{n}-x^{*}\|^{2}-2(1-\alpha_{n})^{2}\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\|g_{n}\|^{2}.$
@@
Next, from Lemma 4.1, there exists $C>0$ such that $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}\leq C$. Moreover, by $L$-Lipschitz continuity of $\nabla f$ and $\nabla f(x^{*})=0$,
$\|\nabla f(x_{n})\|\leq L\|x_{n}-x^{*}\|.$
Hence, by Cauchy–Schwarz,
$\mathbb{E}\big|\langle\nabla f(x_{n}),u-x^{*}\rangle\big|\leq\|u-x^{*}\|\,\mathbb{E}\|\nabla f(x_{n})\|\leq L\|u-x^{*}\|\,\mathbb{E}\|x_{n}-x^{*}\|\leq L\sqrt{C}\,\|u-x^{*}\|.$
Therefore,
$2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\,\mathbb{E}\langle\nabla f(x_{n}),u-x^{*}\rangle\leq 2\alpha_{n}(1-\alpha_{n})\epsilon_{n}L\sqrt{C}\,\|u-x^{*}\|.$
Finally, by the martingale-difference assumption and bounded conditional variance,
@@
$\quad+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|.$ (16)
Set $a_{n}:=\mathbb{E}\|x_{n}-x^{*}\|^{2}$ and define